OpenStudy (luigi0210):

Differentiate:

OpenStudy (luigi0210):

\[\LARGE y=e^{5x^2+7x-13}\]

OpenStudy (agent47):

y' = e^(5x^2+7x-13) * (10x+7)

OpenStudy (agent47):

Simple chain rule man!

OpenStudy (luigi0210):

Oh dang, thank you good sir.

OpenStudy (agent47):

y = e^(f(x)) y' = e^(f(x))* f'(x)

OpenStudy (disco619):

Mam*